OpenStudy (anonymous):

Prove that the following equation is an identity. 1+tan Θ / 1 - tan Θ = cot Θ + 1 / cot Θ - 1

OpenStudy (irishboy123):

you could worse than bust open the tan's and cot's as follows: \[\frac{1+\tan \theta}{ 1 - \tan \theta} • \frac{\cos \theta}{\cos \theta}= \frac{\cot \theta + 1 }{\cot \theta - 1} • \frac{\sin \theta}{\sin \theta}\] and see what you get....